Does the speed of a satellite around the earth depend on its mass? Its distance from the earth? The mass of the earth? Justify your answer.



1
Expert's answer
2021-11-17T10:37:39-0500

"F=ma\to G\frac{Mm}{r^2}=m\frac{v^2}{r}\to v=\sqrt{G\frac{M}{r}}" So,

the speed of a satellite around the Earth depend on its distance from the Earth and the mass of the Earth.